# PhysLLM: Harnessing Large Language Models for Cross-Modal Remote Physiological Sensing

### v1

@@ Abstract
Remote photoplethysmography (rPPG) enables non-contact physiological measurement but remains highly susceptible to illumination changes, motion artifacts, and limited temporal modeling. Large Language Models (LLMs) excel at capturing long-range dependencies, offering a potential solution but struggle with the continuous, noise-sensitive nature of rPPG signals due to their text-centric design. To bridge this gap, we introduce PhysLLM, a collaborative optimization framework that synergizes LLMs with domain-specific rPPG components. Specifically, the Text Prototype Guidance (TPG) strategy is proposed to establish cross-modal alignment by projecting hemodynamic features into LLM-interpretable semantic space, effectively bridging the representational gap between physiological signals and linguistic tokens. Besides, a novel Dual-Domain Stationary (DDS) Algorithm is proposed for resolving signal instability through adaptive time-frequency domain feature re-weighting. Finally, rPPG task-specific cues systematically inject physiological priors through physiological statistics, environmental contextual answering, and task description, leveraging cross-modal learning to integrate both visual and textual information, enabling dynamic adaptation to challenging scenarios like variable illumination and subject movements. Evaluation on four benchmark datasets, PhysLLM achieves state-of-the-art accuracy and robustness, demonstrating superior generalization across lighting variations and motion scenarios.
@@ 1 Introduction
Figure 1: The comparison between (a) typical CNN-based rPPG model, (b) pure LLM model, and (c) our proposed PhysLLM. And (d) is the performance comparison of different methods on the UBFC-rPPG [1] dataset.
Remote photoplethysmography (rPPG) is a non-contact technique that allows for the remote measurement of physiological signals such as heart rate (HR) [43, 3] and blood pressure [36] by analyzing subtle color changes in the skin caused by blood flow. Unlike traditional contact-based methods, such as electrocardiograms (ECG) and Photoplethysmograph (PPG) [20], rPPG does not require physical sensors attached to the body, making it a more convenient and less intrusive option for continuous health monitoring. Based on these advantages of rPPG, more and more excellent work and research have been committed to improving its accuracy, robustness and scalability in recent years.
Traditional remote photoplethysmography (rPPG) methods [23, 19, 4, 4, 34] typically rely on signal processing techniques to isolate rPPG signals from videos. More recent deep learning methods, including CNNs [38, 10, 17, 21, 26, 2, 16] and Transformers [42, 25, 24, 40], have been proposed to address some of these challenges by learning more complex representations of facial features. Fig. 1(a) shows a typical CNN-based framework where features extracted by the encoder are directly fed into the estimator to produce the final sequence. However, these methods remain vulnerable to visual noise (e.g., motion blur, occlusions, low-resolution inputs) and over-reliance on single video streams, limiting their robustness in practical scenarios. The exclusive dependence on visual data creates inherent performance boundaries, particularly in dynamic real-world environments. This underscores the necessity for advanced frameworks that integrate multi-modal inputs and adaptive feature learning through large-scale modeling.
Large Language Models (LLMs) offer significant advantages for enhancing rPPG methods, particularly in modeling long-term temporal dependencies [30, 8]. While traditional rPPG approaches struggle with extended video sequences, LLMs excel at capturing complex sequential patterns, making them promising candidates for physiological signal estimation. Previous works like TimeLLM [8] have demonstrated the effectiveness of LLMs in time-series applications, suggesting untapped potential for rPPG tasks. However, directly applying LLMs to rPPG estimation presents fundamental challenges, as illustrated in Fig. 1(b). The mismatch between LLMs’ discrete operations and rPPG features’ continuous nature leads to poor representations and high noise sensitivity. Despite LLMs’ proven capabilities in cross-modal tasks, their text-oriented architecture requires significant adaptation for effective physiological signal analysis.
@@ 2 Related Work
Remote Physiological Measurement. rPPG has advanced significantly from early signal processing methods to deep learning-based approaches. Traditional techniques, such as blind source separation and skin reflection-based models [23, 19, 4, 34], initially aimed to isolate physiological signals but struggled with real-world variability. Later methods incorporated physiological priors to improve robustness [45], yet they remained sensitive to noise and illumination. With the rise of deep learning, more recent approaches to rPPG measurement have leveraged the power of neural networks to capture rich spatio-temporal representations. Early deep learning-based methods introduced end-to-end spatiotemporal networks [26, 12, 38, 2, 21, 16, 40, 42]. These models demonstrated superior performance compared to traditional methods, as they were capable of capturing intricate temporal and spatial relationships within the video data. One major limitation is their sensitivity to visual interference, such as motion blur, occlusions, and varying lighting conditions, all of which can compromise the quality of the extracted rPPG signal. Recent advancements in vision-language multi-modal learning [44] have shown great promise in enhancing robustness by aligning with additional modalities for rPPG measurement.
LLMs for Time Series Tasks. Recent advancements in LLMs have demonstrated remarkable potential in time series analysis tasks, particularly through their ability to learn temporal patterns and perform cross-modal reasoning. Several pioneering works have explored reprogramming LLMs for time series forecasting without architectural modifications. For instance, Time-LLM [8] introduces a novel reprogramming framework that aligns time series embeddings with LLM token spaces using learnable prompt tokens, achieving state-of-the-art performance across multiple domains. This approach builds upon foundational insights from Time Series Forecasting with LLMs [6], which systematically evaluates LLMs’ inherent temporal reasoning capabilities and proposes specialized fine-tuning strategies to enhance their forecasting accuracy. In healthcare applications, LLMs have shown particular promise for physiological time series interpretation [14]. While existing research primarily focuses on standard physiological signals like ECG and EEG, the application of LLMs to rPPG analysis remains underexplored. Current approaches for rPPG-based health monitoring typically rely on specialized neural architectures, potentially overlooking the cross-modal generalization capabilities inherent in foundation models. Our work bridges this gap by adapting the time series reprogramming module while incorporating rPPG-aware vision-language knowledge, enabling LLMs to interpret subtle cardiovascular patterns from facial videos.
@@ 3.1 Text-Vision-Sequence Embedding Generating
While current deep learning frameworks struggle to extract reliable rPPG signals from facial videos due to environmental interference, existing physiological priors in models like PhysNet [38] demonstrate inherent robustness to such variations. Building on this foundation, we establish a noval architecture with: 1) A fixed PhysNet [38] backbone preserving raw rPPG signals and spatio-temporal features, 2) Three trainable enhancement modules (Dual-Domain Stationary Algorithm module for stable signal tokenization, dedicated multi-scale interaction module for hierarchical feature extraction, Text Prototype Guidanc module aligning these representations with LLM semantic spaces) that progressively refine the physiological representations.
@@
Specifically, let $x\in\mathbb{R}^{B\times L}$ denote the raw rPPG waveform extracted by the PhysNet backbone [38], where batch size $B$ and sequence length $L$ inherit the output dimensions of the backbone. We first compute the global mean $\mu$ and global standard deviation $\sigma$ across the entire sequence,
@@
$\small x^{\prime}=\frac{x-\mu}{\sigma+\epsilon},\quad\epsilon=10^{-5},\vspace{-0.3em}$ (1)
We denote similar standardization processes in a unified manner as function $\mathcal{S}(\cdot)$. To ensure global stationarity, we proceed temporal smoothing with the following operations:
$\small z^{time}_{i}=\alpha\cdot x^{\prime}_{i}+(1-\alpha)\cdot z_{i-1},\quad z_{0}=x^{\prime}_{0},\vspace{-0.2em}$ (2)
where i denotes the current frame value, and $\alpha$ is a smoothing factor. The proof of stationarity is provided in the Appendix. In the following text, we denote these smoothing operations collectively as $\mathcal{Z}(\cdot)$. Therefore, the above operation can be simplified as $z^{time}=\mathcal{Z}(x^{\prime})$.
In parallel, the module performs frequency domain decomposition using the discrete wavelet transform (DWT). The DWT decomposes the input signal into approximation coefficients ($ac$) and detail coefficients ($dc$) on multiple scales. Specifically, the DWT can be expressed as:
$\small x_{ac},[x_{dc,1},\dots,x_{dc,J}]=\text{DWT}(x),\vspace{-0.2em}$ (3)
@@
$\small z^{fre}=\text{IDWT}(\mathcal{Z}(\mathcal{S}(x_{ac})),[\mathcal{Z}(\mathcal{S}(x_{dc,1})),\dots,\mathcal{Z}(\mathcal{S}(x_{dc,J}))]).\vspace{-0.2em}$ (4)
@@
$\small z=(1-\text{$\beta$})\cdot z^{time}+\text{$\beta$}\cdot z^{fre},\vspace{-0.2em}$ (5)
[…]
@@
[…]
@@
[…]
@@
$\small X_{self}=\text{SelfAttention}(\text{$X$}),\vspace{-0.2em}$ (9)
$\small{E^{{}^{\prime}}_{fusion}}=\text{${E^{{}^{\prime}}}+X_{self}$},\vspace{-0.2em}$ (10)
@@
$\small{Output}=\text{${E^{{}^{\prime}}_{fusion}}$ + $y_{2}(E^{{}^{\prime}}_{fusion};X)$},\vspace{-0.2em}$ (12)
$\small\mathcal{T}_{out}=\text{$FFN({E^{{}^{\prime}}_{2}})$},\vspace{-0.2em}$ (13)
where ${E^{{}^{\prime}}_{1}}$ and $E^{{}^{\prime}}_{2}$ are the updated versions of $E^{{}^{\prime}}$. It is worth noting that the TPG module guides not only visual features, but also sequence features, so the input feature X here has the both two meanings. Also, the part of visual features and the part of sequence features shared parameters to learn potential associations.
Table 1: Intra-dataset testing results on UBFC-RPPG [1], PURE [27], BUAA [37] and MMPD [31] datasets. Best results are marked in bold.
@@
ICA [22] 16.00 25.65 0.44 4.77 16.07 0.72 - - - 18.60 24.30 0.01
CHROM [5] 4.06 8.83 0.89 5.77 14.93 0.81 - - - 13.66 18.76 0.08
Green [33] 19.73 31.00 0.37 10.09 23.85 0.34 6.89 10.39 0.60 21.68 27.69 -0.01
POS [35] 4.08 7.72 0.92 3.67 11.82 0.88 - - - 12.36 17.71 0.18
Meta-rPPG [9] 5.97 7.42 0.57 2.52 4.63 0.98 - - - - - -
Contrast-Phys+ [29] 0.21 0.80 0.99 0.48 0.98 0.99 - - - - - -
PhysNet [39] 2.95 3.67 0.97 2.10 2.60 0.99 10.89 11.70 -0.04 4.80 11.80 0.60
PhysFormer [41] 0.92 2.46 0.99 1.10 1.75 0.99 - - - 11.99 18.41 0.18
Efficientphys [13] 1.41 1.81 0.99 4.75 9.39 0.99 16.09 16.80 0.14 13.47 21.32 0.21
@@
3.2 Physiological Cue-Aware Prompt Learning
The extraction of reliable rPPG signals from facial video data presents significant challenges due to varying lighting, subject mobility, and diverse skin tones in complex real-world scenarios. To address these challenges through cross-modal learning, we propose a Physiological Cue-Aware Prompt Learning framework that generates rPPG sample-specific cue and make them guide adaptive token learning.
Cue Caption. Prompt-based adaptation enables LLMs to address specialized subtasks without requiring parameter updates, as demonstrated by the zero-shot capabilities of architectures like LLaMA [32] and DeepSeek [7]. To eliminate manual annotation of context-aware descriptors, we leverage LLaVA [11] for automated physiological cue generation through structured prompt engineering, as shown in Fig. 2(b). For rPPG-specific adaptation, our prompts focus on three critical visual domains: 1) facial anatomical characteristics, 2) transient emotional expressions, and 3) environmental illumination dynamics, as exemplified by our designed query templates in Fig. 4.
Formally, given an input video sequence $V\in\mathbb{R}^{3\times T\times H\times W}$, we extract the central frame $I_{t}$ and process it through LLaVA with task-oriented prompts to obtain detail description, and tokenize the description,
$\small\mathcal{C}_{vision}=Tokenizer(LLaVA(I_{t},Q)),\vspace{-0.2em}$ (14)
where $Q$ denotes our physiological cue queries.
$\small\mathcal{C}_{task}=Tokenizer(\mathcal{P}),\vspace{-0.2em}$ (15)
where $\mathcal{P}$ represents the standardized task preamble.

### v2

@@ Abstract
Remote photoplethysmography (rPPG) enables non-contact physiological measurement but remains highly susceptible to illumination changes, motion artifacts, and limited temporal modeling. Large Language Models (LLMs) excel at capturing long-range dependencies, offering a potential solution but struggle with the continuous, noise-sensitive nature of rPPG signals due to their text-centric design. To bridge this gap, we introduce the PhysLLM, a collaborative optimization framework that synergizes LLMs with domain-specific rPPG components. Specifically, the Text Prototype Guidance (TPG) strategy is proposed to establish cross-modal alignment by projecting hemodynamic features into LLM-interpretable semantic space, effectively bridging the representational gap between physiological signals and linguistic tokens. Besides, a novel Dual-Domain Stationary (DDS) Algorithm is proposed for resolving signal instability through adaptive time-frequency domain feature re-weighting. Finally, rPPG task-specific cues systematically inject physiological priors through physiological statistics, environmental contextual answering, and task description, leveraging cross-modal learning to integrate both visual and textual information, enabling dynamic adaptation to challenging scenarios like variable illumination and subject movements. Evaluation on four benchmark datasets, PhysLLM achieves state-of-the-art accuracy and robustness, demonstrating superior generalization across lighting variations and motion scenarios. The source code is available at https://github.com/Alex036225/PhysLLM.
@@ 1 Introduction
Remote photoplethysmography (rPPG) is a non-contact technique that allows for the remote measurement of physiological signals such as heart rate (HR) (Yue et al., 2021; Das et al., 2023) and blood pressure (Wu et al., 2022) by analyzing subtle color changes in the skin caused by blood flow. Unlike traditional contact-based methods, such as electrocardiograms (ECG) and Photoplethysmograph (PPG) (Murthy et al., 2015), rPPG does not require physical sensors attached to the body, making it a more convenient and less intrusive option for continuous health monitoring. Based on these advantages of rPPG, more and more excellent work and research have been committed to improving its accuracy, robustness and scalability in recent years.
Traditional remote photoplethysmography (rPPG) methods (Poh et al., 2010; Madej et al., 2011; De Haan and Jeanne, 2013; Wang et al., 2017) typically rely on signal processing techniques to isolate rPPG signals from videos. More recent deep learning methods, including CNNs (Yu et al., 2019; Li et al., 2023; Lu et al., 2023; Niu et al., 2019; Špetlík et al., 2018; Chen and McDuff, 2018; Liu et al., 2021b) and Transformers (Yu et al., 2022; Shao et al., 2023; Qian et al., 2024; Yu et al., 2023), have been proposed to address some of these challenges by learning more complex representations of facial features. Fig. 1(a) shows a typical CNN-based framework where features extracted by the encoder are directly fed into the estimator to produce the final sequence. However, these methods are still sensitive to visual noise (e.g., motion blur, occlusions, and low resolution) and often rely on a single video stream, which limits robustness in real-world settings. Introducing textual descriptions provides complementary context about scene conditions—such as occlusion, motion artifacts, and lighting changes—so the model can adapt its processing accordingly. This multi-modal design lets the LLM combine visual cues with semantic context, improving physiological signal extraction under challenging conditions.
Figure 1: The comparison between (a) typical CNN-based rPPG model, (b) pure LLM model, and (c) our proposed PhysLLM.
Large Language Models (LLMs) offer significant advantages for enhancing rPPG methods, particularly in modeling long-term temporal dependencies (Tang et al., 2025; Jin et al., 2023). While traditional rPPG approaches struggle with extended video sequences, LLMs excel at capturing complex sequential patterns, making them promising candidates for physiological signal estimation. Previous works like TimeLLM (Jin et al., 2023) have demonstrated the effectiveness of LLMs in time-series applications, suggesting untapped potential for rPPG tasks. However, directly applying LLMs to rPPG estimation presents fundamental challenges, as illustrated in Fig. 1(b). The mismatch between LLMs’ discrete operations and rPPG features’ continuous nature leads to poor representations and high noise sensitivity. Despite LLMs’ proven capabilities in cross-modal tasks, their text-oriented architecture requires significant adaptation for effective physiological signal analysis.
@@ 2 Related Work
Remote Physiological Measurement. rPPG has advanced significantly from early signal processing methods to deep learning-based approaches. Traditional techniques, such as blind source separation and skin reflection-based models (Poh et al., 2010; Madej et al., 2011; De Haan and Jeanne, 2013; Wang et al., 2017), initially aimed to isolate physiological signals but struggled with real-world variability. Later methods incorporated physiological priors to improve robustness (Zhang et al., 2024), yet they remained sensitive to noise and illumination. With the rise of deep learning, more recent approaches to rPPG measurement have leveraged the power of neural networks to capture rich spatio-temporal representations. Early deep learning-based methods introduced end-to-end spatiotemporal networks (Špetlík et al., 2018; Liu et al., 2020; Yu et al., 2019; Chen and McDuff, 2018; Niu et al., 2019; Liu et al., 2021b; Yu et al., 2023; 2022). These models demonstrated superior performance compared to traditional methods, as they were capable of capturing intricate temporal and spatial relationships within the video data. One major limitation is their sensitivity to visual interference, such as motion blur, occlusions, and varying lighting conditions, all of which can compromise the quality of the extracted rPPG signal. Recent advancements in vision-language multi-modal learning (Yue et al., 2024) have shown great promise in enhancing robustness by aligning with additional modalities for rPPG measurement.
LLMs for Time Series Tasks.  Recent advancements in LLMs have demonstrated remarkable potential in time series analysis tasks, particularly through their ability to learn temporal patterns and perform cross-modal reasoning. Several pioneering works have explored reprogramming LLMs for time series forecasting without architectural modifications. For instance, Time-LLM (Jin et al., 2023) introduces a novel reprogramming framework that aligns time series embeddings with LLM token spaces using learnable prompt tokens, achieving state-of-the-art performance across multiple domains. This approach builds upon foundational insights from Time Series Forecasting with LLMs (Gruver et al., 2023), which systematically evaluates LLMs’ inherent temporal reasoning capabilities and proposes specialized fine-tuning strategies to enhance their forecasting accuracy. In healthcare applications, LLMs have shown particular promise for physiological time series interpretation (Liu et al., 2023b). While existing research primarily focuses on standard physiological signals like ECG and EEG, the application of LLMs to rPPG analysis remains underexplored. Current approaches for rPPG-based health monitoring typically rely on specialized neural architectures, potentially overlooking the cross-modal generalization capabilities inherent in foundation models. Our work bridges this gap by adapting the time series reprogramming module while incorporating rPPG-aware vision-language knowledge, enabling LLMs to interpret subtle cardiovascular patterns from facial videos.
@@ 3.1 Text-Vision-Sequence Embedding Generating
While current deep learning frameworks struggle to extract reliable rPPG signals from facial videos due to environmental interference, existing physiological priors in models like PhysNet (Yu et al., 2019) demonstrate inherent robustness to such variations. Building on this foundation, we establish a novel architecture with: 1) A fixed PhysNet (Yu et al., 2019) backbone preserving raw rPPG signals and spatio-temporal features, 2) Three trainable enhancement modules (Dual-Domain Stationary Algorithm module for stable signal tokenization, dedicated multi-scale interaction module for hierarchical feature extraction, Text Prototype Guidance module aligning these representations with LLM semantic spaces) that progressively refine the physiological representations.
@@
Specifically, let $x\in\mathbb{R}^{B\times L}$ denote the raw rPPG waveform extracted by the PhysNet backbone (Yu et al., 2019), where batch size $B$ and sequence length $L$ inherit the output dimensions of the backbone. We first compute the global mean $\mu$ and global standard deviation $\sigma$ across the entire sequence,
@@
$x^{\prime}=\frac{x-\mu}{\sigma+\epsilon},\quad\epsilon=10^{-5}.\vskip-8.99994pt$ (1)
We denote similar standardization processes in a unified manner as function $\mathcal{H}(\cdot)$. To ensure global stationarity, we proceed temporal smoothing with the following operations:
$z^{time}_{i}=\alpha\cdot x^{\prime}_{i}+(1-\alpha)\cdot z_{i-1},\quad z_{0}=x^{\prime}_{0},\vskip-1.99997pt$ (2)
where i denotes the current frame value, and $\alpha$ is a smoothing factor. The proof of stationarity is provided in Appendix. C. In the following text, we denote these smoothing operations collectively as $\mathcal{Z}(\cdot)$. Therefore, the above operation can be simplified as $z^{time}=\mathcal{Z}(x^{\prime})$.
In parallel, the module performs frequency domain decomposition using the discrete wavelet transform (DWT). The DWT decomposes the input signal into approximation coefficients ($ac$) and detail coefficients ($dc$) on multiple scales. Specifically, the results can be expressed as:
$x_{ac},[x_{dc,1},\dots,x_{dc,J}]=\text{DWT}(x),\vskip-1.99997pt$ (3)
@@
$z^{fre}=\text{IDWT}(\mathcal{Z}(\mathcal{H}(x_{ac})),[\mathcal{Z}(\mathcal{H}(x_{dc,1})),\dots,\mathcal{Z}(\mathcal{H}(x_{dc,J}))]).\vskip-10.00002pt$ (4)
@@
$z=(1-\text{$\beta$})\cdot z^{time}+\text{$\beta$}\cdot z^{fre},\vskip-1.99997pt$ (5)
where $\text{$\beta$}\in[0,1]$ is a learnable parameter.
@@
$\text{$F_{self}$}=\text{SelfAttention}(\text{$F_{cross}$}).\vskip-3.00003pt$ (7)
@@
$\text{$F_{visual}$}=\text{$F_{M}$}+\gamma_{2}\cdot(\text{$F_{cross}$}+\gamma_{1}\cdot\text{$F_{self}$}),\vskip-1.99997pt$ (8)
@@
$X_{self}=\text{SelfAttention}(\text{$X$}),\vskip-3.99994pt$ (9)
${E^{{}^{\prime}}_{fusion}}=\text{${E^{{}^{\prime}}}+X_{self}$},\vskip-3.99994pt$ (10)
$\text{$y_{2}(E^{{}^{\prime}}_{fusion};X)$}=\text{CrossAttention}({E^{{}^{\prime}}_{fusion}},X,X),\vskip-3.99994pt$ (11)
${Output}=\text{${E^{{}^{\prime}}_{fusion}}$ + $y_{2}(E^{{}^{\prime}}_{fusion};X)$},\vskip-1.99997pt$ (12)
$\mathcal{T}_{out}=\text{$FFN({E^{{}^{\prime}}_{2}})$},\vskip-1.99997pt$ (13)
where $E^{{}^{\prime}}_{1}$} and $E^{{}^{\prime}}_{2}$ are the updated versions of $E^{{}^{\prime}}$. It is worth noting that the TPG module guides not only visual features, but also sequence features, so the input feature X here has the two meanings. In addition, the visual feature part and the sequence feature part share the same TPG module to learn potential associations.
3.2 Physiological Cue-Aware Prompt Learning
The extraction of reliable rPPG signals from facial video data presents significant challenges due to varying lighting, subject mobility, and diverse skin tones in complex real-world scenarios. To address these challenges through cross-modal learning, we propose a Physiological Cue-Aware Prompt Learning framework that generates rPPG sample-specific cues for guiding adaptive token learning.
Cue Caption. Prompt-based adaptation enables LLMs to address specialized subtasks without requiring parameter updates, as demonstrated by the zero-shot capabilities of architectures like LLaMA (Touvron et al., 2023) and DeepSeek (Guo et al., 2025). To eliminate manual annotation of context-aware descriptors, we leverage LLaVA (Liu et al., 2023a) for automated physiological cue generation through structured prompt engineering, as shown in Fig. 2(b). For rPPG-specific adaptation, our prompts focus on three critical visual domains: 1) facial anatomical characteristics, 2) transient emotional expressions, and 3) environmental illumination dynamics, as exemplified by our designed query templates in Fig. 4.
Formally, given an input video sequence $V\in\mathbb{R}^{3\times T\times H\times W}$, we extract the central frame $I_{t}$ and process it through LLaVA with task-oriented prompts to obtain detail description, and tokenize the description, $\mathcal{C}_{vision}=Tokenizer(LLaVA(I_{t},Q))$, where $Q$ denotes our physiological cue queries.
$\mathcal{C}_{task}=Tokenizer(\mathcal{P}),\vskip-1.99997pt$ (14)
where $\mathcal{P}$ represents the standardized task preamble.
$\mathcal{S}=\Biggl\{$ $\min_{t}(x_{\mathrm{enc}}[b,t]),\ \max_{t}(x_{\mathrm{enc}}[b,t]),\ \mathrm{median}(x_{\mathrm{enc}}[b,:]),$ (15)
$\Gamma(x_{\mathrm{enc}}[b,:]),\ \mathrm{sgn}\left(\sum_{t=1}^{T-1}(x_{\mathrm{enc}}[b,t+1]-x_{\mathrm{enc}}[b,t])\right),\mathrm{TopK}\left(\mathcal{L}(x_{\mathrm{enc}}[b,:]),5\right)\Biggr\}$
where $\Gamma(\cdot)$ calculates signal trends through first-order differencing, $\Gamma(\mathbf{x})=\sum_{i=2}^{T}(\mathbf{x}_{i}-\mathbf{x}_{i-1}).$
Then $\mathcal{S}$ will make up the symbolic representation as described in Fig. 4(c). Finally, the tokenization process is formulated as $\mathcal{C}_{stats}=\mathrm{Tokenizer}\left(\mathcal{S}\right).$
Adaptive Prompt Learning (APL). Conventional multimodal fusion approaches typically employ static weighting coefficients for prompt integration, suffering from two limitations: (1) fixed combination ratios cannot adapt to learnable LLM, and (2) inability to adaptively select the key cues across different data. As shown in Fig. 2(a), our adaptive fusion paradigm addresses these constraints through learnable hierarchical composition. Formally, having obtained three fundamental cue captions $\mathcal{C}=\{\mathcal{C}_{task},\mathcal{C}_{vision},\mathcal{C}_{stats}\}$, we first implement modality-specific conditioning via dedicated transformation networks,
@@ 3.2 Physiological Cue-Aware Prompt Learning
$\operatorname{AttentiveCompressor}_{k}(x)=\operatorname{Softmax}\left(\frac{Q_{k}K_{k}^{T}}{\sqrt{d}}\right)V_{k},\vskip-3.00003pt$ (17)
with $Q_{k},K_{k},V_{k}$} derived from $x$ through linear projections.
$\mathbf{W}=\left[W_{\text{task}},\ W_{\text{vision}},\ W_{\text{stats}}\right]\in\mathbb{R}^{3\times B\times L\times d},$ (18)
where $L$ is predefined hyperparameter about target prompt sequence length, $d$ is the token embedding dimension, and $W_{\text{task}},W_{\text{vision}},W_{\text{stats}}$ denote learnable tokens, respectively.
The complete fusion operation is formalized as:
$\mathcal{T}_{\text{cue }}=\sum_{\begin{subarray}{c}k\in\Omega\end{subarray}}\mathbf{W}^{(k)}\odot\mathcal{E}_{k},\Omega\triangleq\{task,\\ vision,\\ stats\}.\vskip-3.00003pt$ (19)
Table 1: Intra-dataset testing results on UBFC-RPPG (Bobbia et al., 2019), PURE (Stricker et al., 2014), BUAA (Xi et al., 2020) and MMPD (Tang et al., 2023) datasets. Best results are in bold.
@@
ICA (Poh et al., 2010) 16.00 25.65 0.44 4.77 16.07 0.72 - - - 18.60 24.30 0.01
CHROM (de Haan and Jeanne, 2013) 4.06 8.83 0.89 5.77 14.93 0.81 - - - 13.66 18.76 0.08
Green (Verkruysse et al., 2008) 19.73 31.00 0.37 10.09 23.85 0.34 6.89 10.39 0.60 21.68 27.69 -0.01
POS (Wang et al., 2017) 4.08 7.72 0.92 3.67 11.82 0.88 - - - 12.36 17.71 0.18
Meta-rPPG (Lee et al., 2020) 5.97 7.42 0.57 2.52 4.63 0.98 - - - - - -
PhysNet (Yu et al., 2019) 2.95 3.67 0.97 2.10 2.60 0.99 10.89 11.70 -0.04 4.80 11.80 0.60
PhysFormer (Yu et al., 2022) 0.92 2.46 0.99 1.10 1.75 0.99 8.46 10.17 -0.06 11.99 18.41 0.18
EfficientPhys (Liu et al., 2021a) 1.41 1.81 0.99 4.75 9.39 0.99 16.09 16.80 0.14 13.47 21.32 0.21
Contrast-Phys+ (Sun and Li, 2024) 0.21 0.80 0.99 0.48 0.98 0.99 - - - - - -
RhythmFormer Zou et al. (2025) 0.50 0.78 0.99 0.27 0.47 0.99 9.19 11.93 -0.10 4.69 11.31 0.60
@@
3.3 Training Objectives